OpenStudy (anonymous):

The shadow of a tower when the angle of elevation of the sun is 45o is found to be 10 m longer than when it is 60º. Find the height of the tower

OpenStudy (anonymous):

Then?

OpenStudy (anonymous):

wait. I drew the picture wrong.

OpenStudy (anonymous):



OpenStudy (anonymous):

In the triangle on the RHS, tan(60) = h/x = sqrt(3)

OpenStudy (anonymous):

Now in the big triangle, tan(45) = h/ (x+10)

OpenStudy (anonymous):

This gives two eqns \[x+10 = h\] \[x= \frac{h}{\sqrt{3}}\]

OpenStudy (anonymous):

h= h/ sqrt{3} +10 h [ 1- 1/ sqrt{3}] =10

OpenStudy (anonymous):

\[h= \frac{10}{ 1- \frac{1}{\sqrt{3}}}\]

OpenStudy (anonymous):

h=23.66m